OpenStudy (anonymous):

Express f(x) in the form a(x − h)2 + k. f(x) = x^2 − 6x + 12 what happened..? its answer is f(x) = {(x -3)^2 -9} + 12 =(x-3)^2 + 3 how did this happened..?

OpenStudy (lochana):

f(x) = {(x -3)^2 -9} + 12 =(x-3)^2 + 3

OpenStudy (anonymous):

thats it..? thats the final answer..? u just simplify it..?

OpenStudy (lochana):

yes, now we can say a=1, h=3, k=3

OpenStudy (anonymous):

ahhh.. tnx tnx..

OpenStudy (lochana):

u got it, right?

OpenStudy (anonymous):

yep.. i got it..

OpenStudy (precal):

we used a method called completing the square to convert from standard form to vertex form

OpenStudy (precal):

x^2 − 6x + 12 take the middle number (in this case 6) 6/2 = 3 3^2 is 9 so we are going to +9-9 because this creates a zero

OpenStudy (precal):

x^2-6x+9+12-9 (x-3)^2+12-9 (x-3)^2+3

OpenStudy (precal):

note: (x-3)^2 when you expand it out is x^2-6x+9

OpenStudy (precal):

any other questions

OpenStudy (anonymous):

wow.. sorry.. haha.. tnx for the help.. i now understand how to do it.. tnx..

OpenStudy (precal):

any time here are some quick things to keep in mind always take the middle terms number